# The inverse initial data problem for anisotropic Navier–Stokes equations via Legendre time reduction method

### v2

@@ 6.2 Why use the exponentially weighted time basis
By contrast, using the unweighted Legendre basis $\{Q_{n}\}_{n\geq 0}$ gives $Q_{0}^{\prime}(t)=0$ for all $t$, as $Q_{0}$ is a constant. The corresponding coefficient ${\bf u}_{0}$ can be underrepresented in the derivative equations, which may weaken coupling between modes and degrade numerical conditioning. The exponential weighting avoids this loss by ensuring that even the lowest time mode remains visible to the derivative terms. This observation is confirmed by recomputing Test 1 with the time-dimensional reduction method using the unweighted basis $\{Q_{n}\}_{n\geq 0}$ in $L^{2}(0,T)$ in place of $\{\Psi_{n}\}_{n\geq 0}$; as shown in Figure 4, the exponentially weighted basis $\{\Psi_{n}\}_{n\geq 0}$ yields visibly better reconstructions.
(a) $u_{0}^{{\rm true},1}$
(b) $u_{0}^{{\rm comp},1}$ with the non weighted basis $\{Q_{n}\}_{n\geq 0}$
(c) $u_{0}^{{\rm comp},1}$ with the weighted basis $\{\Psi_{n}\}_{n\geq 0}$, see Figure 1d
(d) $u_{0}^{{\rm true},2}$
(e) $u_{0}^{{\rm comp},2}$ with the non weighted basis $\{Q_{n}\}_{n\geq 0}$
(f) $u_{0}^{{\rm comp},2}$ with the weighted basis $\{\Psi_{n}\}_{n\geq 0}$, see Figure 1e
Figure 4: Test 1: effect of the time basis in the time-dimensional reduction. (a,d) Ground-truth initial fields $u_{0}^{\mathrm{true},1}$ and $u_{0}^{\mathrm{true},2}$. (b,e) Reconstructions using the unweighted Legendre basis $\{Q_{n}\}_{n\geq 0}$ in $L^{2}(0,T)$. (c,f) Reconstructions using the exponentially weighted basis $\{\Psi_{n}\}_{n\geq 0}$ with $\Psi_{n}=e^{t}Q_{n}$ in $L^{2}_{e^{-2t}}(0,T)$. Under identical NN and regularization, the weighted basis yields sharper localization and amplitudes with fewer artifacts than the unweighted alternative.
Figure 4 compares reconstructions in Test 1 using the unweighted Legendre basis $\{Q_{n}\}_{n\geq 0}$ versus the exponentially weighted basis $\{\Psi_{n}\}_{n\geq 0}$ with $\Psi_{n}=e^{t}Q_{n}$ in $L^{2}_{e^{-2t}}(0,T)$. Figures (b,e) show that the unweighted choice produces blurrier shapes and lower peak amplitudes, with noticeable smearing away from the target. In contrast, figures (c,f) obtained with the weighted basis recover crisper boundaries, better peak levels, and fewer artifacts under the same $N$ and regularization. The improvement is consistent with the role of the factor $e^{t}$ in derivative terms: $\Psi_{n}^{\prime}(t)=e^{t}(Q_{n}(t)+Q_{n}^{\prime}(t))$ keeps the lowest time mode visible to $\partial_{t}$ and strengthens the coupling among modes in the time-dimensional reduction. Empirically, this yields a better conditioned inversion and more faithful reconstructions, which supports our choice of the weight function $e^{t}$.
@@ 7 Concluding Remarks
This work presents a new computational approach to the inverse initial data problem for the anisotropic Navier–Stokes equations by leveraging a Legendre-based time reduction technique. By projecting the velocity field onto an exponentially weighted Legendre basis, we reformulated the original time-dependent problem into a sequence of time-independent elliptic systems, making the problem more tractable both analytically and numerically. This dimensional reduction not only simplifies the temporal complexity of the inverse problem but also enables stable recovery of initial velocity fields from partial and noisy boundary measurements.
The numerical results confirm that the proposed method yields accurate and stable reconstructions across diverse test cases, including those with high-contrast and geometrically complex initial conditions. Importantly, the technique exhibits robustness against significant levels of noise, demonstrating its practical potential for real-world applications where direct access to initial states is infeasible.
An important and challenging open question is whether the sequence $\{{\bf U}^{(k)}\}_{k\geq 0}$, defined in Section 4, converges to the true solution of the time-reduced model (3.17). A positive resolution of this question would mark a significant contribution to the theory of inverse problems. One powerful approach to establishing convergence involves incorporating Carleman weight functions into the cost functional $J^{\epsilon,N}_{{\bf U}^{(k)}}$ and utilizing Carleman estimates. However, appropriate Carleman estimates tailored for the anisotropic setting are not currently available.
If we instead restrict our attention to the isotropic case, specifically, the Lamé system, then existing Carleman estimates become applicable, see e.g., [9, 11]. In this context, some recently developed techniques, called the Carleman convexification [12, 14, 15, 19], the Carlaman contraction principle [20, 21, 27] and the Carleman–Newton method [1, 22] can be extended to construct a globally convergent sequence $\{{\bf U}^{(k)}\}_{k\geq 0}$ that approximates the true solution to the time-reduction model (3.17). In this case, both analytical and numerical global solvability of Problem 1 are achievable. Extending these ideas to the anisotropic setting forms a central direction for our future research.

### v4

@@ 6.2 Why use the exponentially weighted time basis
By contrast, using the unweighted Legendre basis $\{Q_{n}\}_{n\geq 0}$ gives $Q_{0}^{\prime}(t)=0$ for all $t$, as $Q_{0}$ is a constant. The corresponding coefficient ${\bf u}_{0}$ can be underrepresented in the derivative equations, which may weaken coupling between modes and degrade numerical conditioning. The exponential weighting avoids this loss by ensuring that even the lowest time mode remains visible to the derivative terms. This observation is confirmed by recomputing Test 3 with the time-dimensional reduction method using the unweighted basis $\{Q_{n}\}_{n\geq 0}$ in $L^{2}(0,T)$ in place of $\{\Psi_{n}\}_{n\geq 0}$; as shown in Figure 4, the exponentially weighted basis $\{\Psi_{n}\}_{n\geq 0}$ yields visibly better reconstructions.
(a) True first velocity component $u^{0}_{\mathrm{true},1}$
(b) Reconstructed first velocity component $u^{0}_{\mathrm{comp},1}$
(c) Relative error of the first component $\frac{|u^{0}_{{\rm true},1}-u^{0}_{{\rm comp},1}|}{\|u^{0}_{{\rm true},1}\|_{L^{\infty}(\Omega)}}$
(d) True second velocity component $u^{0}_{\mathrm{true},2}$
(e) Reconstructed second velocity component $u^{0}_{\mathrm{comp},2}$
(f) Relative error of the second component $\frac{|u^{0}_{{\rm true},2}-u^{0}_{{\rm comp},2}|}{\|u^{0}_{{\rm true},2}\|_{L^{\infty}(\Omega)}}$
Figure 4: Test 3 using the unweighted Legendre basis $\{Q_{n}\}_{n\geq 0}$ in $L^{2}(0,T)$. Figures (a)–(c) show the true first component, its reconstruction, and the corresponding relative error, respectively. Figures (d)–(f) show the true second component, its reconstruction, and the corresponding relative error, respectively.
Figure 4 illustrates the reconstruction in Test 3 when the unweighted Legendre basis $\{Q_{n}\}_{n\geq 0}$ in $L^{2}(0,T)$ is used in place of the exponentially weighted basis $\{\Psi_{n}\}_{n\geq 0}$. Figure 4b and Figure 4e show that, although the main annular and square-ring structures are still identifiable, the reconstructed amplitudes are significantly damped and the interfaces are much more diffuse than those obtained with the weighted basis in Figure 3. In particular, the central disk and the inner square are recovered only in a blurred manner, and noticeable background contamination appears throughout the domain. This degradation is reflected in the relative error distributions shown in Figure 4c and Figure 4f, where the displayed maxima are approximately $0.65$ and $0.67$, respectively, which are substantially larger than those observed in Figure 3. Compared with the results produced by the exponentially weighted basis, the unweighted basis yields weaker contrast, less accurate amplitudes, and poorer resolution of sharp interfaces. These numerical results therefore support the use of the exponentially weighted basis $\{\Psi_{n}\}_{n\geq 0}$ in the time-dimensional reduction procedure.
@@ 7 Concluding Remarks
This work presents a new computational approach to a targeted inverse initial data problem for the compressible anisotropic Navier–Stokes system. The objective is to reconstruct the initial velocity field ${\bf u}^{0}$ from noisy boundary observations, while the density $\rho({\bf x},t)$, the pressure $p({\bf x},t)$, the anisotropic viscosity tensor $\bm{\mu}$, and the body force ${\bf F}({\bf x},t)$ are assumed known. The proposed approach is based on an exponentially weighted Legendre time-dimensional reduction, which transforms the original time-dependent inverse problem into a coupled system of time-independent elliptic equations for the Fourier coefficients of the velocity field. This reduction makes the problem substantially more tractable both analytically and computationally.
To solve the resulting time-reduced model, we combine quasi-reversibility with a damped Picard iteration. The numerical results demonstrate that the proposed method yields accurate and stable reconstructions across several test cases, including examples with localized structures, non-axis-aligned features, multiple disconnected regions, and sign changes. The experiments also show robustness with respect to noisy boundary data, which indicates that the method has promising potential for inverse problems in realistic settings where the initial state is not directly observable.
An important open question is whether the sequence $\bigl\{\mathbf{U}^{(k)}\bigr\}_{k\geq 0}$ defined in Section 4 converges to the true solution of the time-reduced model (3.15). A positive answer would constitute a significant advance in the analysis of the proposed method. One possible route toward such a result would be to incorporate Carleman weights into the functional $J_{{\bf U}^{(k)}}^{\epsilon,N}$ and to combine this construction with suitable Carleman estimates. However, such estimates are not currently available in the fully anisotropic setting considered here.
If one restricts instead to the isotropic case, in particular to the Lamé system, then existing Carleman estimates become applicable; see, e.g., [9, 11]. In that setting, several recently developed techniques, including Carleman convexification [12, 14, 15, 19], the Carleman contraction principle [20, 21, 27, 29], and the Carleman–Newton method [1, 22], may be adapted to construct a globally convergent sequence $\bigl\{\mathbf{U}^{(k)}\bigr\}_{k\geq 0}$ that approximates the true solution of the time-reduced model (3.15). In such a case, both analytical and numerical global solvability of Problem 1 may become achievable. Extending these ideas to the anisotropic setting remains a central direction for future research.